### v1

# Compatibility of canonical $\ell$-adic local systems on Shimura varieties

Abstract.
Let $(G,X)$ be a Shimura datum. In previous work [16] with Klevdal, we showed that the canonical $G(\mathbb{Q}_{\ell})$-valued local systems on Shimura varieties for $G$ form compatible systems after projection to the adjoint group of $G$. In this note, we strengthen this result to prove compatibility for the $G(\mathbb{Q}_{\ell})$-local systems themselves. We also include the crystalline compatibility, extending the adjoint case established in our joint work [9] with Huryn, Kedlaya, and Klevdal.
I thank Christian Klevdal and Jake Huryn for discussions related to this material and for their comments on an earlier version.
1. Introduction
Let $(G,X)$ be a Shimura datum, and let $K_{0}\subset G(\mathbb{A}_{f})$ be a neat compact open subgroup. The associated Shimura variety $\mathrm{Sh}_{K_{0}}:=\mathrm{Sh}_{K_{0}}(G,X)$ is a smooth quasi-projective variety over the reflex field $E(G,X)$, with underlying complex manifold $\mathrm{Sh}_{K_{0}}(G,X)(\mathbb{C})=G(\mathbb{Q})\backslash(X\times G(\mathbb{A}_{f}))/K_{0}$. Varying the compact open subgroup $K\subset K_{0}$ we have a morphism of varieties over $E(G,X)$, $\mathrm{Sh}_{K}\to\mathrm{Sh}_{K_{0}}$, and we write $\mathrm{Sh}=\lim_{K}\mathrm{Sh}_{K}$. Fix a basepoint $s\in\mathrm{Sh}(\mathbb{C})$, inducing basepoints in all finite-level Shimura varieties $\mathrm{Sh}_{K}$ as well. We then obtain the geometrically connected component $S_{K,s}$ of $\mathrm{Sh}_{K}$ containing $s$; it is defined over some finite (abelian) extension $E_{K,s}$ of $E(G,X)$. When $Z_{G}(\mathbb{Q})\subset Z_{G}(\mathbb{A}_{f})$ is a discrete subgroup, for any normal compact open subgroup $K\subset K_{0}$, $\mathrm{Sh}_{K}\to\mathrm{Sh}_{K_{0}}$ is a Galois cover with Galois group $K_{0}/K$, and (via the choice of $s$) these assemble to canonical local systems
$\pi_{1}(S_{K_{0},s},s)\to K_{0},$
which for varying $\ell$ give us by projection the canonical $\ell$-adic local systems
$\rho_{\ell}\colon\pi_{1}(S_{K_{0},s},s)\to G(\mathbb{Q}_{\ell}).$
A long-standing hope ([6]) is that $\mathrm{Sh}_{K_{0}}$ parametrizes a family of motives; this gives rise to the expectation that the $\rho_{\ell}$ form a compatible system of $\ell$-adic representations, at least in the following sense:
  1. (1)
    There is an integer $N$ (independent of $\ell$), and an integral model $\mathcal{S}_{K_{0},s}$ over $\mathcal{O}_{E_{K_{0},s}}[1/N]$, such that for all $\ell$, $\rho_{\ell}$ extends to (factors through) an arithmetic local system
    $\rho_{\ell}\colon\pi_{1}(\mathcal{S}_{K_{0},s},s)\to G(\mathbb{Q}_{\ell}).$
  2. (2)
    For all $\ell$, for all closed points $x$ of (the finite-type $\mathbb{Z}$-scheme) $\mathcal{S}_{K_{0},s}[1/\ell]$, the semisimple conjugacy class underlying $\rho_{\ell}(\mathrm{Frob}_{x})$ is defined over $\overline{\mathbb{Q}}$ and is independent of $\ell$.
For more refined expectations, see the end of this introduction. Much is known about problems (1) and (2). When $(G,X)$ is of abelian type, these and much more are known by work of Kisin ([12], [13]). When $(G,X)$ is not of abelian type—let us for simplicity at present assume that $G^{\mathrm{ad}}$ is $\mathbb{Q}$-simple—then $\mathrm{rk}_{\mathbb{R}}(G^{\mathrm{ad}}_{\mathbb{R}})\geq 2$, and our joint work [16] with Klevdal proved that (1) and (2) hold after projecting to the adjoint local systems $\rho_{\ell}^{\mathrm{ad}}\colon\pi_{1}(S_{K_{0},s},s)\to G^{\mathrm{ad}}(\mathbb{Q}_{\ell})$.
The present paper resolves (1) and (2) in general, establishing also some refinements in this generality: $\mathbb{Q}$-rationality of the independent of $\ell$-conjugacy classes in (2), and crystalline compatibility, which builds on the analogous result in the adjoint case in our joint work [9] with Huryn, Kedlaya, and Klevdal.
Theorem 1.1.
Let $(G,X)$ be a Shimura datum such that $Z_{G}(\mathbb{Q})$ is a discrete subgroup of $Z_{G}(\mathbb{A}_{f})$, let $K_{0}\subset G(\mathbb{A}_{f})$ be a neat compact open subgroup, and let $s\in\mathrm{Sh}(\mathbb{C})$, $S_{K_{0},s}$ be as above. Assume that for all $\mathbb{Q}$-simple factors $H$ of $G^{\mathrm{ad}}$, $\mathrm{rk}_{\mathbb{R}}(H_{\mathbb{R}})\geq 2$. Then there is an integer $N$ and an integral model $\mathcal{S}:=\mathcal{S}_{K_{0},s}$ over $\mathcal{O}_{E_{K_{0},s}}[1/N]$ such that
  1. (1)
    For all $\ell$, $\rho_{\ell}$ factors through
    $\pi_{1}(\mathcal{S},s)\to G(\mathbb{Q}_{\ell}).$
  2. (2)
    For all closed points $x\in\mathcal{S}_{K_{0},s}[1/\ell]$, the class of $\rho_{\ell}(\mathrm{Frob}_{x})$ in $[G{\,\!\sslash\!\,}G](\overline{\mathbb{Q}}_{\ell})$ lies in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ and is independent of $\ell$ (not equal to the residue characteristic of $x$). [^1] In fact, $\rho_{\ell}(\mathrm{Frob}_{x})$ is semisimple, by [3, Theorem 7.1]
  3. (3)
    When $p$ is the residue characteristic of the closed point $x$, which lies in the special fiber $\mathcal{S}_{\kappa(v)}$ for some $v|p$, the local system $\rho_{p}$ is crystalline, has an associated overconvergent $G$-valued $F$-isocrystal $\rho_{p,v}^{\operatorname{F-Isoc}^{\dagger}}$ on $\mathcal{S}_{\kappa(v)}$, and the linearized crystalline Frobenius of $\rho_{p,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$ defines the same class in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ as $\rho_{\ell}(\mathrm{Frob}_{x})$ for $\ell\neq p$.
We have stated this theorem with the real rank condition on factors of $G^{\mathrm{ad}}$ to be clear what can be deduced using our previous work [16]; but of course, combined with Kisin’s work in abelian type, which is needed to handle real rank 1 cases, this yields:
Corollary 1.2.
Let $(G,X)$ be a Shimura datum such that $Z_{G}(\mathbb{Q})$ is a discrete subgroup of $Z_{G}(\mathbb{A}_{f})$, let $K_{0}\subset G(\mathbb{A}_{f})$ be a neat compact open subgroup, and let $s\in\mathrm{Sh}(\mathbb{C})$, $S_{K_{0},s}$ be as above. Then all the conclusions of Theorem 1.1 hold.
Finally, we remark that the “pointwise” compatibility statements in these results are really consequences of the stronger statements that for fixed $v$, all $\rho_{\ell}|_{\mathcal{S}_{\kappa(v)}}$ (and $\rho_{p,v}^{\operatorname{F-Isoc}^{\dagger}}$) are $G$-companions in the strong sense of [7]. The proofs will clarify this assertion.
Theorem 1.1 results from combining our previous work [16] with the work of Bakker-Shankar-Tsimerman ([3]) showing that all Shimura varieties have canonical integral models away from a finite set of bad primes (essentially the $N$ of Theorem 1.1). We should clarify the logical relation between these three papers: [3] makes use of Part (1) of Theorem 1.1 (see Lemma 5.1, Theorem 5.2 of loc. cit.), which we established for $\rho_{\ell}^{\mathrm{ad}}$ in [16, Theorem 2.3]. In §3, we give a proof of Part (1) of Theorem 1.1 in general that does not depend on the two earlier papers. Using their main result on integral models along with our main theorem from [16] (the analogue of Part (2) of Theorem 1.1 for $\rho_{\ell}^{\mathrm{ad}}$), Bakker-Shankar Tsimerman prove ([3, Theorem 9.7]) that for $v\nmid N$, every $\mu$-ordinary point of the special fiber $\mathcal{S}_{v}$ lifts to a characteristic zero special point. [^2] To be precise, [3] use our earlier result for $\rho_{\ell}^{\mathrm{ad}}$ to cover the non-abelian type (or the real rank at most 1) cases; in abelian type, [13, Corollary 2.3.1] establishes the desired compatibility. We in turn use this result on CM lifting of $\mu$-ordinary points to establish Parts (2) and (3) of Theorem 1.1.
We now summarize each section and the arguments therein. In §2, we recall results of Drinfeld ([7]) on Deligne’s companions conjecture; we need his work on companions for representations valued in not necessarily connected reductive (not only semisimple) groups, and so we elaborate slightly on the results in [7, §6]. In §3, we show (Corollary 3.4) that Part (1) of Theorem 1.1 holds: there is an integer $N$ such that for all $\ell$, all $\rho_{\ell}$ are unramified away from $N\ell$. For the $G^{\mathrm{ad}}$-representations, we showed this in [16, Proposition 2.3] by an abstract argument using superrigidity and Deligne’s conjecture. When $G$ is not an adjoint group, descents to $\pi_{1}(S_{K_{0},s})$ of a given representation of $\pi_{1}^{\mathrm{top}}(S_{K_{0},s}(\mathbb{C}))$ (extended to $\pi_{1}((S_{K_{0},s})_{\overline{\mathbb{Q}}})$ using integrality) are not unique, and it is easy to choose descents for varying $\ell$ that have no common finite bad set $N$ as in Part (1) of the Theorem. We need to use the fact that our descents arise from the family of canonical models; it turns out that all we need in addition to the arguments in the adjoint case is that there is a number field point $y$ of $S_{K_{0},s}$ for which the specialized local systems $\rho_{\ell}|_{y}$ are unramified outside $N_{y}\ell$ for some integer $N_{y}$ independent of $\ell$. In §4 we prove Parts (2) and (3) of Theorem 1.1. Given the result of §3, from [3, Theorem 9.2] we deduce that on the $\mu$-ordinary locus of the special fiber $\mathcal{S}_{v}$ ($v\nmid N\ell\ell^{\prime}$), $\rho_{\ell}$ and $\rho_{\ell^{\prime}}$ satisfy the companions property (2). Using the companions construction (§2) and Zariski-density of the $\mu$-ordinary locus, we can deduce the companions property on all of $\mathcal{S}_{v}$ via a strong multiplicity 1 property for Galois representations. The crystalline compatibility works analogously but requires as additional input the overconvergent ($G$-valued) F-isocrystals associated to the crystalline (by Esnault-Groechenig: [19, Appendix]) local systems $\rho_{p}$ in our joint work [9].
We end this introduction by remarking on what we do not do, and what we hope to pursue in future work. We view our results in the framework of the Langlands-Rapoport conjecture, as a significant step toward associating Kottwitz triples to mod $p$ points on Shimura varieties; in abelian type, Kisin has proven the Langlands-Rapoport conjecture ([13]), but this remains a wide-open and fundamental problem in non-abelian type. We further note that Kottwitz triples are not only a useful intermediary in work on the full Langlands-Rapoport conjecture (as in loc. cit.) but are in principle a sufficient intermediary to calculate the zeta function of the Shimura variety; indeed, this is how Kottwitz first employed them in [14]. Recall ([13, §4.3]) that a Kottwitz triple over $k=\mathbb{F}_{p^{r}}$ is a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$ where
  • $\gamma_{0}\in G(\mathbb{Q})$, well-defined up to $G(\overline{\mathbb{Q}})$-conjugacy;
  • $(\gamma_{\ell})_{\ell\neq p}\in G(\mathbb{A}^{p}_{f})$;
  • $\delta\in G(\mathrm{Frac}(W(k)))$, well-defined up to Frobenius conjugacy by elements of $G(W(k))$.
These data are required to satisfy several conditions:
  • $\gamma_{0}$ is $G(\overline{\mathbb{A}}^{p}_{f})$-conjugate to $(\gamma_{\ell})_{\ell\neq p}$, where $\overline{\mathbb{A}}_{f}^{p}$ is the restricted direct product of the $\overline{\mathbb{Q}}_{\ell}$ ($\ell\neq p$) with respect to the $\overline{\mathbb{Z}}_{\ell}$.
  • $\gamma_{0}$ is $G(\overline{\mathbb{Q}}_{p})$-conjugate to $\gamma_{p}:=\delta\sigma(\delta)\cdots\sigma^{r-1}(\delta)$.
  • $\gamma_{0}\in G(\mathbb{R})$ is elliptic.
  • After possibly replacing $k$ by a finite extension, there is an inner twisting $I$ of $I_{0}:=\mathrm{Cent}_{G}(\gamma_{0})$ such that $I\otimes_{\mathbb{Q}}\mathbb{R}$ is anisotropic mod center, and for all primes $\ell$, $I\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ is isomorphic to $I_{\ell}$ as an inner twist of $I_{0}\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ (here $I_{\ell}$ for $\ell\neq p$ is $\mathrm{Cent}_{G_{\mathbb{Q}_{\ell}}}(\gamma_{\ell})$, and for $\ell=p$ it is the Frobenius-centralizer: see [13, §4.3.1] for details).
To every $k$-point $x\in\mathcal{S}(k)$ ($p\nmid N$) we associate a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$, except that we have only constructed $\gamma_{0}$ in general as an element of $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. With that caveat, we have verified the first and second desiderata for the triple: $\gamma_{\ell}$ is of course $\rho_{\ell}(\mathrm{Frob}_{x})$, and $\delta$ arises from the absolute Frobenius ($\gamma_{p}$ being the linearized version) on the $F$-isocrystal at $x$ arising from the crystalline local system $\rho_{p}$; Kisin’s integral $p$-adic Hodge theory ([12, §2]) gives the precise well-definedness property of $\delta$. The key question remaining seems to be (for $x$ not in the $\mu$-ordinary locus, where there is no issue) finding the actual rational representative $\gamma\in G(\mathbb{Q})$; we hope to return to this problem (and the verification of the remaining conditions) in future work.
2. Recollection on companions
For a smooth variety $X$ over a finite field $k\supset\mathbb{F}_{p}$ and a geometric point $\xi$ of $X$, set $\pi_{1}(X)=\pi_{1}(X,\xi)$. Drinfeld proves ([7, Theorem 1.4.1]) that for any semisimple, not necessarily connected, group $H$ over $\overline{\mathbb{Q}}$, any prime $\lambda$ not above $p$ of $\overline{\mathbb{Q}}$, and any continuous representation $\rho_{\lambda}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda})$ with Zariski-dense image, $\rho_{\lambda}$ has for any other place $\lambda^{\prime}\nmid p$ a companion
$\rho_{\lambda\leadsto\lambda^{\prime}}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda^{\prime}}),$
also with Zariski-dense image. This $\rho_{\lambda\leadsto\lambda^{\prime}}$ has the weak companion property that for all closed points $x\in X$, $\rho_{\lambda}(\mathrm{Frob}_{x})$ and $\rho_{\lambda\leadsto\lambda^{\prime}}(\mathrm{Frob}_{x})$ define a common element of the GIT quotient (space of conjugacy classes) $[H{\,\!\sslash\!\,}H](\overline{\mathbb{Q}})$ (inside the ¯$\overline{\mathbb{Q}}_{\lambda}$ and ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$ points, respectively). It also has a deeper strong companion property that characterizes it uniquely even in the absence of the implication “local Frobenius conjugacy implies global conjugacy.” Namely, for fixed $\lambda$, let $\hat{\Pi}_{\lambda}$ be the inverse limit over the targets of all Zariski-dense continuous homomorphisms $\pi_{1}(X,\xi)\to H(\overline{\mathbb{Q}}_{\lambda})$. Let $\widetilde{X}$ be a “universal cover” of $X$, i.e. a pro-object of finite étale covers that pro-represents a fiber functor on $\mathrm{F\acute{E}t}(X)$, let $\Pi=\operatorname{Aut}(\widetilde{X}/X)$, and choose a geometric point $\tilde{\xi}$ of $\widetilde{X}$ over our geometric point $\xi$. This choice induces an isomorphism $\Pi\cong\pi_{1}(X,\xi)^{\mathrm{op}}$, or (composing with $g\mapsto g^{-1}$) if one prefers an isomorphism $\Pi\cong\pi_{1}(X,\xi)$. $\Pi$ contains a canonical subset $\Pi_{\mathrm{Frob}}$ of Frobenius elements $\{\mathrm{Frob}_{\tilde{x}}^{n}:\tilde{x}\in|\widetilde{X}|,n\geq 1\}$ coming from the classical theory of decomposition groups, and our chosen isomorphism transfers this subset to $\pi_{1}(X,\xi)$. The groups $\hat{\Pi}_{\lambda}$ descend, uniquely up to inner automorphisms by the identity component, to groups $\hat{\Pi}_{(\lambda)}$ over $\overline{\mathbb{Q}}$. Drinfeld shows that for two places $\lambda$, $\lambda^{\prime}$, there is a unique isomorphism $\hat{\Pi}_{(\lambda)}\cong\hat{\Pi}_{(\lambda^{\prime})}$, modulo inner automorphisms by the identity component, preserving the canonical diagrams
$\Pi_{\mathrm{Frob}}\to[\hat{\Pi}_{(\lambda)}{\,\!\sslash\!\,}\hat{\Pi}^{0}_{(\lambda)}](\overline{\mathbb{Q}})\to\Pi=\pi_{0}(\hat{\Pi}_{(\lambda)}).$
This isomorphism produces the companions $\rho_{\lambda\to\lambda^{\prime}}$. We will not make explicit use of this stronger companion property, but we will note when (as will be the case) our assertions of compatibility have this stronger meaning.
In the present paper we will need an essentially equivalent variant with reductive monodromy groups. Drinfeld presents one such in [7, §6], where he defines a group $\Pi^{\mathrm{mot}}$ over $\overline{\mathbb{Q}}$ that similarly allows for the construction of weak or strong companions for continuous homomorphisms $\rho_{\lambda}\colon\pi_{(}X)\to H(\overline{\mathbb{Q}}_{\lambda})$ with reductive (not necessarily connected) algebraic monodromy group such that in each representation of $H$ and for all $x\in|X|$, the eigenvalues of $\mathrm{Frob}_{x}$ are all #$\#\kappa(x)$-Weil numbers. We explain another variant here.
Lemma 2.1.
Let $\rho_{\lambda}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda})$ be a continuous homomorphism with Zariski-dense image in a reductive (not necessarily connected) group $H/\overline{\mathbb{Q}}$. Assume that $\rho_{\lambda}$ is plain of characteristic $p$ in the following sense (c.f. [4, §4]): for all (finite-dimensional algebraic) representations $r$ of $H$ and all $x\in|X|$, the eigenvalues of $r(\rho_{\lambda}(\mathrm{Frob}_{x}))$ lie in ¯¯$\overline{\mathbb{Q}}\subset\overline{\mathbb{Q}}_{\lambda}$ and are $v$-adic units for all finite places $v$ not dividing $p$. Then $\rho_{\lambda}$ has a strong $\lambda^{\prime}$-companion for any finite place $\lambda^{\prime}$ of $\overline{\mathbb{Q}}$.
Proof.
The proof will explain the meaning of the final conclusion. Following [7, §6], let $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ be the pro-reductive $\lambda$-adic completion of $\Pi$. By Proposition 3.5.2 of loc. cit., the canonical map
$\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}\times_{\widehat{\mathbb{Z}}}\mathrm{Hom}(\overline{\mathbb{Z}}_{\lambda}^{\times},\mathbb{G}_{m})$
is an isomorphism. (The two maps are the canonical projection $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}$ and the map induced by
¯$\overline{\mathbb{Z}}_{\lambda}^{\times}\xrightarrow{\sim}\mathrm{Hom}_{\mathrm{cts}}(\widehat{\mathbb{Z}},\overline{\mathbb{Q}}_{\lambda}^{\times})\to\mathrm{Hom}_{\mathrm{cts}}(\pi_{1}(X),\overline{\mathbb{Q}}_{\lambda}^{\times})=\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m}),$
the middle arrow being pullback along $\pi_{1}(X)\to\mathrm{Gal}(\bar{k}/\mathbb{F}_{p})\simeq\widehat{\mathbb{Z}}$.) Replacing ¯$\overline{\mathbb{Z}}_{\lambda}^{\times}$ by its subgroup $\mathcal{W}_{p}\subset\overline{\mathbb{Q}}^{\times}\cap\overline{\mathbb{Z}}_{\lambda}^{\times}$ (intersection inside ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$) of all $p$-Weil numbers, Drinfeld obtains a pro-reductive group scheme quotient $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ whose finite-dimensional representations are those of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ such that the corresponding $\lambda$-adic representation has, for all $x\in|X|$, $\mathrm{Frob}_{x}$-eigenvalues that are $q_{x}$-Weil numbers. Exactly the same construction applies if instead of $\mathcal{W}_{p}$ we take the subgroup
$\mathcal{P}_{p}\subset\overline{\mathbb{Q}}^{\times}\cap\overline{\mathbb{Z}}_{\lambda}^{\times}$
of algebraic numbers that are plain of characteristic $p$ (units at all finite places away from $p$), yielding an intermediate quotient
$\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{plain}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{mot}}$
classifying continuous semisimple $\lambda$-adic representations whose Frobenius eigenvalues are all plain of characteristic $p$ [^3] We give a few details not made so explicit in [7, §6]. Either the Weil number or plainness conclusions can be checked after replacing $X$ by a finite étale cover. Thus for a given representation $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow H$ (which in the two cases we would assume factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ or $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$), we can assume it factors through $H^{\mathrm{geo},0}\cdot Z_{H^{0}}^{0}$, with $H^{\mathrm{geo}}$ denoting the geometric monodromy group. Projecting, we get a representation $\hat{\Pi}^{\mathrm{red}}_{\lambda}\to H^{\mathrm{geo},0}/H^{\mathrm{geo},0}\cap Z_{H^{0}}^{0}$ that factors, by semisimplicity of geometric monodromy, through $\hat{\Pi}_{\lambda}$; again replacing $X$ by a finite étale cover, we may assume this lifts to a homomorphism $\hat{\Pi}_{\lambda}\to H^{\mathrm{geo},0}$. Similarly projecting to $Z^{0}_{H^{0}}/H^{\mathrm{geo},0}\cap Z_{H^{0}}^{0}$ and replacing $X$ by a finite étale cover that allows us to lift to $Z^{0}_{H^{0}}$, we get a homomorphism $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to Z^{0}_{H^{0}}$ that factors through $\mathrm{Hom}(A,\mathbb{G}_{m})$, with $A=\overline{\mathbb{Z}}_{\lambda}^{\times}$, $\mathcal{W}_{p}$, or $\mathcal{P}_{p}$ according to our assumptions on the original representation. (Note that $\ker(\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\mathrm{Hom}(A,\mathbb{G}_{m}))$ is isomorphic to $\ker(\hat{\Pi}_{\lambda}\to\hat{\mathbb{Z}})$ and factors through $H^{\mathrm{geo},0}$ in our $H$-representation.) In summary, after replacing $X$ by a finite cover, we can assume our homomorphism is a product of a $\hat{\Pi}_{\lambda}\to H^{\mathrm{geo},0}$ and a $\mathrm{Hom}(A,\mathbb{G}_{m})\to Z^{0}_{H^{0}}\cong\mathbb{G}_{m}^{r}$, i.e. an $r$-tuple $(\alpha_{1},\ldots,\alpha_{r})\in A^{r}$. In any finite-dimensional representation of $H$, the former sends any $\mathrm{Frob}_{x}$ to a matrix with pure weight 0 eigenvalues, and the latter sends $\mathrm{Frob}_{x}$ to a product $\prod_{i=1}^{r}\alpha_{i}^{d_{i}\deg(x)}$, where the $d_{i}$ are integers depending on the representation and $\deg(x)=[\kappa(x):\mathbb{F}_{p}]$. Thus representations of $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ and $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ give $\lambda$-adic representations with Weil number and plain eigenvalues as claimed. The converse is similar. $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$ are both independent of $\lambda$, so using the main theorem [7, Theorem 1.4.1] on $\lambda$-independence of $\hat{\Pi}_{\lambda}$ we can define groups over $\overline{\mathbb{Q}}$, $\hat{\Pi}^{\mathrm{plain}}=\hat{\Pi}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$ (and likewise $\hat{\Pi}^{\mathrm{mot}}$, as already done in [7, §6.1]) such that for all, $\lambda$ $\hat{\Pi}^{\mathrm{plain}}\times_{\overline{\mathbb{Q}}}\overline{\mathbb{Q}}_{\lambda}$ is isomorphic, uniquely up to conjugation by $\hat{\Pi}_{\lambda}^{\mathrm{plain},0}$, to $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$.
In particular, in the setting of the Lemma, we have a surjection $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{plain}}\twoheadrightarrow H_{\overline{\mathbb{Q}}_{\lambda}}$. Descending the second arrow to $\overline{\mathbb{Q}}$, we get $\hat{\Pi}^{\mathrm{plain}}\twoheadrightarrow H$ (unique up to $H^{0}$-conjugation), which in turn for any $\lambda^{\prime}$ extends to $\hat{\Pi}_{\lambda^{\prime}}^{\mathrm{plain}}\twoheadrightarrow H_{\overline{\mathbb{Q}}_{\lambda^{\prime}}}$, which classifies the companion $\rho_{\lambda\leadsto\lambda^{\prime}}$.
3. Ramification
In [16, Proposition 2.3], we showed that Zariski-dense $G^{\mathrm{ad}}(\mathbb{C})$-valued representations $\rho$ of $\pi_{1}^{\mathrm{top}}(X(\mathbb{C}))$ in the “superrigid” (loc. cit. Definition 2.2) regime have arithmetic descents with independent of $\ell$ ramification (the same would work for Zariski-dense cohomologically rigid representations with quasi-unipotent local monodromy and finite-order abelianization). Without Zariski-density, and more precisely when the centralizer of $\rho$ is non-trivial, the argument breaks down, and indeed there is no reason that arbitrarily chosen arithmetic descents $\rho_{\lambda}\colon\pi_{1}(X_{F})\to G(\overline{\mathbb{Q}}_{\lambda})$ should have this property, since one can for varying $\lambda$ twist $\rho_{\lambda}$ by homomorphisms $\mathrm{Gal}(\overline{\mathbb{Q}}/F)\to Z_{G}(\rho_{\lambda})(\overline{\mathbb{Q}}_{\lambda})$ with different ramification sets (the basic case to keep in mind, even with $\rho^{\mathrm{ad}}$ Zariski-dense, is when this centralizer is the center of a non-adjoint group $G$). In this section (Lemma 3.2), we show that having one number field specialization with independent of $\ell$ ramification is enough to show that a collection of descents also has this property.
Let $G$ be a connected reductive group; since we will work with sufficiently large coefficients, we assume $G$ is split over $\mathbb{Z}$. In this section we work with a smooth quasi-projective variety $X$ over a number field $F$ whose topological fundamental group $\Gamma=\pi_{1}^{\mathrm{top}}(X(\mathbb{C}),\bar{x})$ (we will specify a base-point later) is $G^{\mathrm{ad}}$-superrigid in the sense of [16, Definition 2.2]: that is, for any algebraically closed field $\Omega$ of characteristic zero and any two Zariski-dense homomorphisms $\rho_{1},\rho_{2}\colon\Gamma\to G^{\mathrm{ad}}(\Omega)$, there is a $\tau\in\operatorname{Aut}(G^{\mathrm{ad}}_{\Omega})$ such that $\tau(\rho_{1})=\rho_{2}$. [^4] Alternatively, throughout this section we could take cohomologically rigid representations with bounded order of quasi-unipotency, without assuming superrigidity properties of $\Gamma$; somewhat longer versions of the arguments then apply using the technique of [8], [15]. We fix once and for all a smooth projective compactification $X\subset\overline{X}$ with $\overline{X}\setminus X$ a strict normal crossings divisor. We further let $N$ be a large enough integer that the setup $(X,\overline{X},\overline{X}\setminus X)$ spreads out to a relative strict normal crossings divisor over $\mathcal{O}_{F}[1/N]$, and we choose some $x\in X(\mathcal{O}_{F^{\prime}}[1/N])$ with $F^{\prime}/F$ a finite extension.
For a fixed finite subgroup $Z\subset Z_{G}$, which we may assume contains $Z_{G^{\mathrm{der}}}$, let $\mathcal{R}_{Z}$ be the set of $G(\mathbb{C})$-conjugacy classes of homomorphisms $\rho\colon\Gamma\to G^{\mathrm{der}}Z(\mathbb{C})$ such that $\rho^{\mathrm{ad}}$ is Zariski-dense in $G^{\mathrm{ad}}$ (fixing $Z$ amounts to bounding the abelianization). Set $G_{Z}=G^{\mathrm{der}}\cdot Z$. Since $\Gamma$ is $G^{\mathrm{ad}}$-superrigid, $\mathcal{R}_{Z}$ is finite. As in [8], [15], all members of $\mathcal{R}_{Z}$ are integral, so we may choose representatives such that each $\rho\in\mathcal{R}_{Z}$ factors through $G(\mathcal{O}_{L})$ for some number field $L\subset\overline{\mathbb{Q}}$. In turn, for each place $\lambda$ of $\overline{\mathbb{Q}}$ we obtain the representations
$\rho_{\lambda,\overline{\mathbb{Q}}}\colon\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})\to G_{Z}(\overline{\mathbb{Q}}_{\lambda}).$
Lemma 3.1.
There is an integer $N^{\prime}$ such that for every $\rho\in\mathcal{R}_{Z}$, for every $\lambda$, and for every place $\bar{v}$ of $\overline{\mathbb{Q}}$ not dividing $N^{\prime}\ell$, $\rho_{\lambda,\overline{\mathbb{Q}}}$ factors through the tame specialization map $\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})\xrightarrow{\mathrm{sp}}\pi_{1}^{t}(X_{\kappa(\bar{v})},\bar{x}_{v})$.
Proof.
For each $\lambda$, there is a finite extension $F(\lambda)/F$ such that for all $\rho\in\mathcal{R}_{Z}$, $\rho_{\lambda,\overline{\mathbb{Q}}}$ descends to a continuous homomorphism
$\rho_{\lambda,Z}\colon\pi_{1}(X_{F(\lambda)},\bar{x})\to G_{Z}(\overline{\mathbb{Q}}_{\lambda}).$
Indeed, this follows from a standard argument (variant of [21, Theorem 4], [8], [15]): the outer action of $\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ on $\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})$ induces a discrete action on the finite set $\mathcal{R}_{Z}$, so there is an open subgroup $H\subset\mathrm{Gal}(\overline{\mathbb{Q}}/F)$ such that $H$ fixes each $\rho_{\lambda,\overline{\mathbb{Q}}}$ (up to isomorphism). The adjoint descent $\rho_{\lambda}^{\mathrm{ad}}$ exists uniquely (using Zariski-density: see [16, Lemma 2.1]) over $F(\lambda)^{\mathrm{ad}}:=\overline{\mathbb{Q}}^{H}$. As we vary $\rho$, the $\rho_{\lambda}^{\mathrm{ad}}$ all factor through $G^{\mathrm{ad}}(\mathcal{O})$ for $\mathcal{O}$ the ring of integers in a finite extension of $\mathbb{Q}_{\ell}$, and $G_{Z}(\mathcal{O})\to G^{\mathrm{ad}}(\mathcal{O})$ is on an open subgroup of the source an isomorphism onto its image (an open subgroup of the target). Therefore after restricting to a finite extension $F(\lambda)/F(\lambda)^{\mathrm{ad}}$, each $\rho_{\lambda}^{\mathrm{ad}}$ lifts to
$\rho_{\lambda}\colon\pi_{1}(X_{F(\lambda)},\bar{x})\to G_{Z}(\mathcal{O}).$
Each $\rho_{\lambda}$ factors through $\pi_{1}(X_{\mathcal{O}_{F(\lambda)}[1/N_{\lambda}]},\bar{x})$ for an integer $N_{\lambda}$ (which we may assume divisible by $\ell$) that depends on $\lambda$ but not on $\rho\in\mathcal{R}_{Z}$. As in [16, Proposition 2.3], set $N^{\prime}=\mathrm{gcd}(N_{\lambda})$, and suppose that $p$ does not divide $N^{\prime}$. We claim that for all $\lambda^{\prime}\nmid p$, $\rho_{\lambda^{\prime},\overline{\mathbb{Q}}}$ factors via the tame specialization map through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ for every $v|p$. Suppose that $p$ divides $N_{\lambda^{\prime}}$ (else the assertion is clear). For some $\lambda\neq\lambda^{\prime}$, $p$ does not divide $N_{\lambda}$. Let $\bar{v}$ be any place of $\overline{\mathbb{Q}}$ above $p$, and let $v=\bar{v}|_{F(\lambda)}$, so for each $\rho\in\mathcal{R}_{Z}$ we have the (tame: see [16, Footnote 3]) restriction to the special fiber
$\rho_{\lambda,v}\colon\pi_{1}(X_{\kappa(v)},\bar{x}_{v})\to G_{Z}(\overline{\mathbb{Q}}_{\lambda}).$
By [7, Theorem 1.4.1], $\rho_{\lambda,v}$ has a (strong) $\lambda^{\prime}$-companion
$\rho_{\lambda\leadsto\lambda^{\prime},v}\colon\pi_{1}(X_{\kappa(v)},\bar{x}_{v})\to G_{Z}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$
with the “same” algebraic monodromy group as $\rho_{\lambda,v}$. (We don’t assert this is all of $G_{Z}$; it is some group between $G^{\mathrm{der}}$ and $G_{Z}$.) Since $\rho_{\lambda\leadsto\lambda^{\prime},v}$ is also tamely ramified, we can as in [8], [15] pull back along the tame specialization map to a homomorphism
$\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}\colon\pi_{1}^{\mathrm{top}}(X(\mathbb{C}),\bar{x})\to G_{Z}(\overline{\mathbb{Q}}_{\lambda^{\prime}}),$
which clearly lies in $\mathcal{R}_{Z}$. [^6] Had we not assumed superrigidity and instead assumed the elements of $\mathcal{R}_{Z}$ to be cohomologically rigid, this conclusion would still be valid but would require the arguments of [8] and [15]. We may view the elements of $\mathcal{R}_{Z}$ as ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$-valued using that they all have representatives factoring through $G(\mathcal{O}_{L})\subset G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$, and we claim that the map $\mathcal{R}_{Z}\to\mathcal{R}_{Z}$, $\rho\mapsto\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$, is a bijection. Indeed, let $\rho$ and $\sigma$ be elements of $\mathcal{R}_{Z}$ such that $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}^{\mathrm{top}}$ (equality up to $G$-conjugacy). Then $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime}\bar{v}}$. Let $F_{v}\in\pi_{1}(X_{\kappa(v)},\bar{x}_{v})$ be the $v$-Frobenius induced by the section ($\kappa(v)$-rational point) $x_{v}$. The equality $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}(F_{v}\gamma F_{v}^{-1})$, for all $\gamma\in\pi_{1}(X_{\overline{\kappa(v)}},\bar{x}_{v})$ implies that $\sigma_{\lambda\leadsto\lambda^{\prime},v}(F_{v})^{-1}\rho_{\lambda\leadsto\lambda^{\prime},v}(F_{v})$ centralizes the image of $\rho_{\lambda\leadsto\lambda^{\prime},\bar{v}}=\sigma_{\lambda\leadsto\lambda^{\prime},\bar{v}}$, hence lies in $Z$. As $Z$ is finite and central, passing to a finite extension $\kappa^{\prime}/\kappa(v)$, we see $\sigma_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda\leadsto\lambda^{\prime},v}|_{X_{\kappa^{\prime}}}$. Applying [7, Theorem 1.4.1] again shows that $\sigma_{\lambda,v}|_{X_{\kappa^{\prime}}}=\rho_{\lambda,v}|_{X_{\kappa^{\prime}}}$, so $\sigma_{\lambda,\bar{v}}=\rho_{\lambda,\bar{v}}$, and finally $\sigma=\rho$.
Lemma 3.2.
Assume that for some $\rho\in\mathcal{R}_{Z}$, there are for all $\lambda$ descents $\rho_{\lambda}$ as above such that
  1. (1)
    The fields $F(\lambda)$ are independent of $\lambda$.
  2. (2)
    For our given rational point $x\in X(\mathcal{O}_{F^{\prime}}[1/N])$, there is an integer $N_{x}$ (divisible by $N$) such that the specialization $\rho_{\lambda,x}$ is unramified outside $N_{x}\ell$ for all places $\lambda$.
Then there is an integer $N^{\prime\prime}$ such that for every $\lambda$, $\rho_{\lambda}$ factors through $\pi_{1}(X_{\mathcal{O}_{F}[1/N^{\prime\prime}\ell]},\bar{x})$.
Proof.
For notational simplicity we assume $F(\lambda)=F$ for all $\lambda$ (if this entails enlarging $F$, we can correspondingly enlarge $F^{\prime}$). The point $x$ induces a splitting
$\pi_{1}(X_{F^{\prime}},\bar{x})\simeq\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})\rtimes\mathrm{Gal}(\overline{\mathbb{Q}}/F^{\prime}).$
Let $N^{\prime\prime}$ be the product of the integer $N^{\prime}$ from Lemma 3.1, the primes (below those places) ramified in $F^{\prime}/F$, and the integer $N_{x}$. Let $v|p$ be a place of $F$ such that $v$ does not divide $N^{\prime\prime}\ell$. Let $w$ be a place of $F^{\prime}$ above $v$. Let $\mathcal{O}^{\mathrm{h}}$ be the henselization of the local ring $\mathcal{O}_{F^{\prime},{(w)}}$, let $F^{\mathrm{h}}$ be $\mathrm{Frac}(\mathcal{O}^{\mathrm{h}})$, let $\mathcal{O}^{\mathrm{sh}}$ the strict henselization of $\mathcal{O}_{F^{\prime},{(w)}}$, and let $F^{\mathrm{sh}}=\mathrm{Frac}(\mathcal{O}^{\mathrm{sh}})$, all constructed as subrings of $\overline{\mathbb{Q}}$. [^7] We hope the notation $\mathcal{O}^{\mathrm{h}}$, etc., causes no confusion, since it does not refer to $F^{\prime}$ and $w$; it instead has the advantage of being uncluttered. I think one should actually work over $\mathcal{O}$ the henselization of the local ring $\mathcal{O}_{F^{\prime},(w)}$. The analysis below applies there, and this connects more easily to the global theory: to see a finite map with normal source $X^{\prime}\to X_{\mathcal{O}_{F}^{\prime}}[1/N]$ is étale above . We will show that $\rho_{\lambda}|_{X_{F^{\mathrm{h}}}}$ factors through $\pi_{1}(X_{\mathcal{O}^{\mathrm{h}}})$. Restricting the splitting $x$ to
$\pi_{1}(X_{F^{\mathrm{h}}},\bar{x})\simeq\pi_{1}(X_{\overline{\mathbb{Q}}},\bar{x})\rtimes\mathrm{Gal}(\overline{F}/F^{\mathrm{h}}),$
we obtain a factorization
${\pi_{1}(X_{F^{\mathrm{h}}},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{\lambda}}$${\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})\rtimes\pi_{1}(\mathrm{Spec}(\mathcal{O}^{\mathrm{h}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G(\overline{\mathbb{Q}}_{\lambda}).}$ (1)
Indeed, by Lemma 3.1 $\rho_{\lambda,\overline{\mathbb{Q}}}$ factors through $\pi_{1}^{t}(X_{\overline{\kappa(v)}},\bar{x})\cong\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})$ (this isomorphism follows from tame specialization: see [17, Corollary A.12] and [16, Footnote 3]), and $\rho_{\lambda,x}|_{F^{\mathrm{h}}}$ is unramified, giving the two components of the factorization. The restriction of $\rho_{\lambda}$ to $X_{F^{h}}$ therefore factors through $\pi_{1}(X_{\mathcal{O}^{\mathrm{h}}},\bar{x})$, as this group is isomorphic to $\pi_{1}(X_{\mathcal{O}^{\mathrm{sh}}},\bar{x})\rtimes\pi_{1}(\mathrm{Spec}(\mathcal{O}^{\mathrm{h}}))$, compatibly with the canonical maps from $\pi_{1}(X_{F^{\mathrm{h}}},\bar{x})$: this last point follows by Lemma 3.3, an analogue over a henselian DVR of the exact sequence linking arithmetic and geometric fundamental groups. We are done. [^8] Details: For each $r$, let $Y_{r}\to X[1/N_{\lambda}]$ be the finite étale covering corresponding to the mod $\lambda^{r}$ reduction over a suitable finite extension of $\mathbb{Q}_{\ell}$, and let $\overline{Y}_{r}$ be the normalization of $X^{\prime}=X[1/N_{\lambda}]\cup X_{w}$ in the function field of $Y_{r}$. We must check $\overline{Y}_{r}\to X^{\prime}$ is étale in the fiber above $w$. This is clearly equivalent to checking $\overline{Y}_{r}\times\mathcal{O}_{(w)}\to X^{\prime}_{\mathcal{O}_{(w)}}$ is étale. We are given that $\overline{Y}_{r}\times\mathcal{O}^{\mathrm{h}}\to X^{\prime}_{\mathcal{O}^{\mathrm{h}}}$ is étale. Thus for some finite étale (since $\mathcal{O}_{(w)}$ is a DVR, étale neighborhoods of the special point are explicit) $\mathcal{O}_{(w)}\to\mathcal{O}^{\prime}$, $\overline{Y}_{r}\times\mathcal{O}^{\prime}\to X^{\prime}_{\mathcal{O}^{\prime}}$ is étale (by a limit argument). Thus $\overline{Y}_{r}\to X^{\prime}$ is étale by étale descent.
Lemma 3.3.
Let $\mathcal{O}$ be a henselian DVR, and let $X\to\mathrm{Spec}(\mathcal{O})$ be a quasi-projective scheme such that $X_{\mathcal{O}^{\mathrm{sh}}}$ is connected. Let $\bar{x}\in X(\overline{K})$ be a geometric point with $K=\mathrm{Frac}(\mathcal{O})$. Then the canonical maps induce a short exact sequence
[…]
Proof.
We imitate [22, Lemma 0BTX]. Surjectivity of $\beta$ is clear: connected objects of $\mathrm{F\acute{E}t}(\mathcal{O})$ are all isomorphic to finite étale subextensions $\mathcal{O}\to\mathcal{O}^{\prime}\subset\mathcal{O}^{\mathrm{sh}}$, and $X_{\mathcal{O}^{\prime}}$ is then connected by assumption. The composite $\beta\circ\alpha$ is trivial; note that for $\mathcal{O}^{\prime}$ as before, $\mathcal{O^{\prime}}\otimes_{\mathcal{O}}\mathcal{O}^{\mathrm{sh}}\equiv\prod_{\mathrm{Gal}(K^{\prime}/K)}\mathcal{O}^{\mathrm{sh}}$ (with $K^{\prime}=\mathrm{Frac}(\mathcal{O}^{\prime})$, $K=\mathrm{Frac}(\mathcal{O})$), just as for fields. To show the image of $\alpha$ is normal, we check that for $U\to X$ finite étale with $U$ connected and such that $U_{\mathcal{O}^{\mathrm{sh}}}\to X_{\mathcal{O}^{\mathrm{sh}}}$ has a section $s$, $U_{\mathcal{O}^{\mathrm{sh}}}$ is a finite coproduct of copies of $X_{\mathcal{O}^{\mathrm{sh}}}$. There is an $\mathcal{O}^{\prime}$ as above such that $s$ is defined over $\mathcal{O}^{\prime}$, and $s(X_{\mathcal{O}^{\prime}})$ is an open connected component of $U_{\mathcal{O}^{\prime}}$ (being a section of an étale separated morphism). Now consider
$V^{\prime}=\bigcup_{\sigma\in\mathrm{Gal}(K^{\prime}/K)}s^{\sigma}(X_{\mathcal{O}^{\prime}})\subset U_{\mathcal{O}^{\prime}}.$
$V^{\prime}$ has a canonical $\mathrm{Gal}(K^{\prime}/K)$-action compatible with that on $\mathcal{O}^{\prime}$, and $\mathcal{O}\to\mathcal{O}^{\prime}$ is a Galois cover with group $\mathrm{Gal}(K^{\prime}/K)$, so by Galois descent there is an open connected component $V$ of $U$ descending $V^{\prime}$. Since $U$ is connected, $V=U$, and thus $U_{\mathcal{O}^{\mathrm{sh}}}=V_{\mathcal{O}_{\mathrm{sh}}}$ is a disjoint union of copies of $X_{\mathcal{O}^{\mathrm{sh}}}$. Injectivity of $\alpha$: any finite étale $V\to X_{\mathcal{O}^{\mathrm{sh}}}$ arises from $V^{\prime}\to X_{\mathcal{O}^{\prime}}$ for some finite étale $\mathcal{O}\to\mathcal{O}^{\prime}$ as above. Then $V^{\prime}\to X$ is still finite étale, and $V^{\prime}\times_{\mathcal{O}}\mathcal{O}^{\mathrm{sh}}$ contains $V=V^{\prime}\times_{\mathcal{O}^{\prime}}\mathcal{O}^{\mathrm{sh}}$ as an open and closed subscheme (again using the Galois property $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\mathcal{O}^{\mathrm{sh}}\cong\prod_{\mathrm{Gal}(K^{\prime}/K)}\mathcal{O}^{\mathrm{sh}})$. Finally, to show exactness in the middle it remains to check that for any finite étale $U\to X$ such that $U_{\mathcal{O}^{\mathrm{sh}}}$ is a finite disjoint union of copies of $X_{\mathcal{O}^{\mathrm{sh}}}$, there is a finite étale $S\to\mathrm{Spec}(\mathcal{O})$ and a surjection $X\times_{\mathcal{O}}S\to U$. We first find a finite étale $\mathcal{O}\to\mathcal{O}^{\prime}$ as above such that $U_{\mathcal{O}^{\prime}}=\sqcup_{i=1}^{n}X_{\mathcal{O^{\prime}}}$. $S=\sqcup_{i=1}^{n}\mathrm{Spec}(\mathcal{O}^{\prime})$ works. ∎
Corollary 3.4.
For a Shimura datum $(G,X)$ such that $Z_{G}(\mathbb{Q})$ is a discrete subgroup of $Z_{G}(\mathbb{A}_{f})$, a neat level structure $K_{0}$, and a basepoint $s\in\mathrm{Sh}(\mathbb{C})$, there is an integer $N$ such that for all $\ell$ the canonical local system
$\rho_{\ell}\colon\pi_{1}(S_{K_{0},s},\bar{s})\to G(\mathbb{Q}_{\ell})$
is unramified outside $N\ell$, i.e. extends to $\mathcal{S}_{K_{0},s}[1/\ell]$ for an integral model $\mathcal{S}_{K_{0},s}$ over $\mathcal{O}_{E_{K_{0},s}}[1/N]$.
In particular, this completes the proof in the non-adjoint case of [3, Lemma 5.1], specifically the part deduced from Theorem 5.2 of loc. cit..
4. Shimura Varieties
Set $\mathcal{S}=\mathcal{S}_{K_{0},s}$, our integral model over $\mathcal{O}_{E_{K_{0},s}}[1/N]$. Here $N$ is chosen large enough that the conclusion of Corollary 3.4 holds for $N$, and that [3, Theorem 1.3] holds away from this $N$, i.e. for $v\nmid N$, $\mathcal{S}_{K_{0}}\times_{\mathcal{O}_{E}[1/N]}\mathcal{O}_{E_{v}}$ is an integral canonical model. We have the canonical local systems
$\rho_{\ell}\colon\pi_{1}(\mathcal{S}[1/\ell])\to G(\mathbb{Q}_{\ell}),$
which we will show in this section form a compatible system. For places $\lambda|\ell$, $\lambda^{\prime}|\ell^{\prime}$ of $\overline{\mathbb{Q}}$, we also denote $\rho_{\ell}$ (resp. $\rho_{\ell^{\prime}}$) by $\rho_{\lambda}$ (resp. $\rho_{\lambda^{\prime}}$) when viewed as valued in $G(\overline{\mathbb{Q}}_{\lambda})$ (resp. $G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. Let $v$ be a prime of $E_{K_{0},s}$ not dividing $N\ell\ell^{\prime}$. We form $\rho_{\lambda,v}$, the restriction to the $v$-fiber. Since the monodromy groups of the $\rho_{\lambda,v}$ might not be semisimple, we need a small argument to check that Drinfeld’s results ([7]) apply. As explained in Lemma 2.1, for the results of [7, §6] to apply to give a $\lambda^{\prime}$-companion, it suffices to check that the eigenvalues of $r(\rho_{\lambda,v})(\mathrm{Frob}_{x})$ are plain of characteristic $p$ for all $x\in|\mathcal{S}_{\kappa(v)}|$ and representations $r$ of the algebraic monodromy group $G_{\lambda,v}\subset G$ of $\rho_{\lambda,v}$. To check this, it suffices to check that the composite $\pi_{1}(\mathcal{S}_{v})\to G/G^{\mathrm{der}}$ is plain of characteristic $p$, but this follows from the canonical model property on the zero-dimensional Shimura variety associated to $G/G^{\mathrm{der}}$, and in particular is a special case of Lemma 4.1 below. Thus a strong $\lambda^{\prime}$-adic companion $\rho_{\lambda\leadsto\lambda^{\prime},v}$ exists.
Lemma 4.1.
Let $s=[x,a]\in\mathrm{Sh}(\mathbb{C})$ be a special point defined in $S_{K_{0},s}$ over a number field $E(sK_{0})$. Then the canonical local system $\rho_{K_{0},s}$ specialized to $sK_{0}$ defines a compatible system of $G(\mathbb{Q}_{\ell})$-representations in the following sense: for any choice of Frobenius $\mathrm{Frob}_{v}$ at a place $v|p$ ($p$ not dividing $N$ as above) of $E(sK_{0})$, $\rho_{\ell,s}(\mathrm{Frob}_{v})\in G(\mathbb{Q}_{\ell})$ is for all $\ell\neq p$ $G(\mathbb{Q}_{\ell})$-conjugate to an independent of $\ell$ value $q_{\mathrm{Frob}_{v}}^{-1}\in G(\mathbb{Q})$. When $\ell=p$, as we vary the finite-dimensional representation $\xi$ of $G$, the F-isocrystal $D_{\mathrm{cris}}(\xi\circ\rho_{p,s}|_{\Gamma_{E(sK_{0})_{v}}})$ on $\mathrm{Spec}(\kappa(v))$ with its linearized crystalline Frobenius $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ defines up to conjugacy an element $\gamma_{v}\in G(\overline{\mathbb{Q}}_{p})$, and $\gamma_{v}=q_{\mathrm{Frob}_{v}}^{-1}$ in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$.
Proof.
This is a standard calculation, but we give some details particularly since we have no reference for the crystalline compatibility. Let $\rho_{s}$ be the specialization of $\rho_{K_{0}}$ along $s$ and likewise $\rho_{\ell,s}$ for its $G(\mathbb{Q}_{\ell})$ projection for all $\ell$. By [16, Lemma 3.2], for any $\alpha\in\mathbb{A}_{E(sK_{0})}^{\times}$ mapping to $\sigma\in\Gamma^{\mathrm{ab}}_{E(sK_{0})}$, we have
$q_{\sigma}a\rho_{s}(\sigma)=r_{x}(\alpha)a,$
i.e., $\rho_{s}(\sigma)=a^{-1}q_{\sigma}^{-1}r_{x}(\alpha)a$, for some $q_{\sigma}\in G(\mathbb{Q})_{+}$. For $\sigma\in\Gamma_{E(sK_{0})_{v}}$ and $\ell\neq p$, by $\ell$-projection we obtain $\rho_{\ell,s}(\sigma)=a_{\ell}^{-1}q_{\sigma}^{-1}a_{\ell}$; since the special point is ¯$\overline{\mathbb{Z}}[1/N]$-rational, for $\sigma\in I_{E(sK_{0})_{v}}$ we get $q_{\sigma}=1$, and for $\sigma=\mathrm{Frob}_{v}$ we find $\rho_{\ell,s}(\mathrm{Frob}_{v})$ is for all $\ell\neq p$ conjugate (in $G(\mathbb{Q}_{\ell})$) to the independent of $\ell$ value $q_{\mathrm{Frob}_{v}}^{-1}$.
Projecting to the $p$-component, for $\sigma\in I_{E(sK_{0})_{v}}$ we thus find $\rho_{p,s}(\sigma)=a_{p}^{-1}r_{x}(\alpha)a_{p}$, where $\alpha\in\mathcal{O}_{v}^{\times}$, and $\rho_{p,s}(\mathrm{Frob}_{v})=a_{p}^{-1}q_{\mathrm{Frob}_{v}}^{-1}r_{x}(\varpi_{v})a_{p}$ for the uniformizer $\varpi_{v}$ mapped to the choice of $\mathrm{Frob}_{v}$. In particular, up to conjugacy the abelian representation $\rho_{p,s}|_{I_{E(sK_{0})_{v}}}$ is given by the algebraic representation $r_{x}$, hence is crystalline. In particular, the functor $\mathrm{Rep}(G_{\mathbb{Q}_{p}})\to\operatorname{F-Isoc}(\kappa(v))$,
$\xi\mapsto D_{\mathrm{cris}}(\xi\circ\rho_{p,s}),$
composed with the canonical (forgetful) fiber functor to $\mathrm{Vect}_{E(sK_{0})_{v}}$, gives by the Tannakian formalism an element $\gamma_{v}\in G(\overline{\mathbb{Q}}_{p})$, well-defined up to inner automorphism (resulting from choosing an isomorphism over some finite extension of $E(sK_{0})_{v}$ between this fiber functor and the canonical fiber functor on $\mathrm{Rep}(G_{\mathbb{Q}_{p}})$), such that for all $\xi\in\mathrm{Rep}(G_{\mathbb{Q}_{p}})$, $\xi(\gamma_{v})$ has the same characteristic polynomial (and even same conjugacy class) as the linearized crystalline Frobenius $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ acting on $D_{\mathrm{cris}}(\xi\circ\rho_{p,s})$. We check this is the characteristic polynomial of $\xi(\operatorname{inn}(a_{p})(\rho_{p,s}(\mathrm{Frob}_{v}))\cdot r_{x}(\varpi_{v})^{-1})$, i.e. is that of $\xi(q_{\mathrm{Frob}_{v}}^{-1})$.
There is a finite extension $K/\mathbb{Q}_{p}$ such that for any representation $\xi$ of $G_{\mathbb{Q}_{p}}$,
$\xi_{K}\circ\operatorname{inn}(a_{p})(\rho_{p,s}|_{\Gamma_{E(sK_{0})_{v}}})=\bigoplus_{i}\psi_{i}$
for crystalline characters $\psi_{i}\colon\Gamma_{E(sK_{0})_{v}}\to K^{\times}$. Restricting to $I_{E(sK_{0})_{v}}$, this decomposition induces a decomposition $\xi_{K}\circ r_{x}=\bigoplus_{i}\chi_{i}$ for algebraic homomorphisms $\chi_{i}\colon\operatorname{Res}_{E(sK_{0})_{v}/\mathbb{Q}_{p}}\mathbb{G}_{m}\to\operatorname{Res}_{K/\mathbb{Q}_{p}}\mathbb{G}_{m}$ of $\mathbb{Q}_{p}$-tori. A calculation that ultimately rests on Lubin-Tate theory ([5, Proposition B.4], to which strictly speaking we should add the corresponding calculation of crystalline periods for unramified representations) shows that $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ acts on $D_{\mathrm{cris}}(\xi_{K}\circ\operatorname{inn}(a_{p})(\rho_{p,s}|_{\Gamma_{E(sK_{0})_{v}}}))$ (canonically a direct sum of free rank 1 $E(sK_{0})_{v}\otimes_{\mathbb{Q}_{p}}K$-modules, with $\varphi^{[\kappa(v):\mathbb{Q}_{p}]}$ acting $E(sK_{0})_{v}$-linearly) by
$\bigoplus_{i}\psi_{i}(\operatorname{rec}_{v}(\varpi_{v}))\chi_{i}(\varpi_{v})^{-1}.$
By the calculation of $\rho_{p,s}$, this implies $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ has the same characteristic polynomial as
$\xi_{K}\left(\operatorname{inn}(a_{p})(\rho_{p,s}(\varpi_{v}))\cdot r_{x}(\varpi_{v})^{-1}\right)=\xi_{K}(q_{\mathrm{Frob}_{v}}^{-1}).$
and thus the same characteristic polynomial as $\xi(q_{\mathrm{Frob}_{v}}^{-1})$.
Since finite-dimensional representations separate semi-simple conjugacy classes, $\rho_{\ell,s}(\mathrm{Frob}_{v})$, $q_{\mathrm{Frob}_{v}^{-1}}$, and $\gamma_{v}$ are equal as elements of $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. ∎
To show $\rho_{\lambda\leadsto\lambda^{\prime},v}$ is conjugate to $\rho_{\lambda^{\prime},v}$, we will use results of [3] that depend on the adjoint case of Theorem 4.4, established in [16]. We begin by reviewing these.
Theorem 4.2 (Theorem 1.6 of [3]).
Let $v$ be a prime of $E_{K_{0},s}$ not dividing $N$. For any $x\in\mathcal{S}_{K_{0}}(\mathbb{F}_{q})$ lying in the $\mu$-ordinary locus, there is a (canonical) special point $\tilde{x}\in\mathcal{S}_{K_{0}}(W(\mathbb{F}_{q}))$ lifting $x$.
[…]
Corollary 4.3.
Let $v|p$ be a prime of $E_{K_{0},s}$ not dividing $N$. For any closed point $x\in\mathcal{S}_{\kappa(v)}$ belonging to the $\mu$-ordinary locus and all $\ell\neq p$, $\rho_{\ell}(\mathrm{Frob}_{x})$ is $G(\mathbb{Q}_{\ell})$-conjugate to an $-\ell$ element of $G(\mathbb{Q})$.
Theorem 4.4.
For all finite places $\lambda$ and $\lambda^{\prime}$ of $\overline{\mathbb{Q}}$ and all $v$ not dividing $N\ell\ell^{\prime}$, $\rho_{\lambda^{\prime},v}=\rho_{\lambda\leadsto\lambda^{\prime},v}$ as $G^{\mathrm{der}}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$-conjugacy classes of homomorphisms $\pi_{1}(\mathcal{S}_{\kappa(v)})\to G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$.
Proof.
Let $H$ be the algebraic monodromy group of $\rho_{\lambda^{\prime},v}$. Our setup implies that $H$ lies between $G^{\mathrm{der}}$ and $G$ (but may be neither semisimple nor connected). From the known adjoint case ([16, Theorem 3.10]) and the abelian compatibility (Lemma 4.1), after suitable conjugation we may assume $\rho_{\lambda\leadsto\lambda^{\prime},v}=\rho_{\lambda^{\prime},v}\cdot\chi$ for some character (depending on $\lambda$, $\lambda^{\prime}$, and $v$) $\chi\colon\pi_{1}(\mathcal{S}_{\kappa(v)})\to Z_{G^{\mathrm{der}}}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. We know from Theorem 4.2 that $\rho_{\lambda\leadsto\lambda^{\prime},v}(\mathrm{Frob}_{x})$ and $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ are conjugate for closed points $x$ in the Dirichlet density 1 (because Zariski open and dense: see Lemma 4.5) $\mu$-ordinary locus of $\mathcal{S}_{\kappa(v)}$. In particular, for any irreducible representation $R\colon G\to\mathrm{GL}_{N}$, by the Čebotarev density theorem (and Brauer-Nesbitt theorem), there is $g\in\mathrm{GL}_{N}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$ such that
$g(R\circ\rho_{\lambda^{\prime},v})g^{-1}=R\circ\rho_{\lambda\leadsto\lambda^{\prime},v}=R\circ\rho_{\lambda^{\prime},v}\cdot R\circ\chi.$
In particular, $\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))=\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))\cdot R(\chi)$, where here we view $R(\chi)$ as a scalar (by Schur’s lemma) in ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}^{\times}$. Now, either $R(\chi)(\mathrm{Frob}_{x})=1$ for a density 1 set of closed points $x$, in which case $R(\chi)=1$, or there is a positive density set of $x$ such that $\mathrm{tr}(R(\rho_{\lambda^{\prime},v})(\mathrm{Frob}_{x}))$ equals zero. In the latter case, by [20, Theorem 3], there is a connected component $\Phi$ of the algebraic monodromy group $H_{R}$ of $R(\rho_{\lambda^{\prime},v})$ on which the trace vanishes identically. [^10] Compare [18, Proposition 3.4.9]. Strictly speaking, Rajan proves his Theorem for representations of $\pi_{1}(\mathcal{O}_{F}[1/N])$, $F$ a number field, but the same arguments will apply for integral finite type $\mathbb{Z}$-schemes, where we still have the Čebotarev density theorem. For some $z\in Z_{G}$, $\Phi=R(zH^{0})$ ($H$ surjects onto $H_{R}$, $H^{0}$ surjects onto $H_{R}^{0}$, and every element of $H/H^{0}$ is represented by an element of $Z_{G}$ since $H\supset G^{\mathrm{der}}$), and since $R(z)$ is a scalar, $\mathrm{tr}(\Phi)=0$ forces $\mathrm{tr}(R(H^{0}))=0$; this is clearly impossible, since $1\in H^{0}$. We conclude that $R(\chi)=1$. Knowing this for all irreducible representations $R$ of $G$, we conclude that $\chi=1$.
Here is the simple lemma cited in the course of the proof of Theorem 4.4:
Lemma 4.5.
Let $X$ be a scheme of finite-type over $\mathbb{Z}$. Let $U\subset X$ be a Zariski open and dense subset. Then the closed points $|U|$ have Dirichlet density 1 in $X$.
Proof.
For all such $X$, $\zeta_{X}(s)=\sum_{x\in|X|}(|\kappa(x)|)^{-s}$ converges absolutely and uniformly in $\Re(s)>d$ but has a pole at $s=d$ (the former statement is elementary, while the latter uses at least the Riemann hypothesis for curves over finite fields). If $U\subset X$ is open and Zariski-dense, then $Z=X\setminus U$ has dimension less than $d$, so $\zeta_{Z}(s)$ converges in $\Re(s)>d-1$. The density of $|U|$ is by definition
$\lim_{s\to d^{+}}\frac{\sum_{x\in|U|}(|\kappa(x)|)^{-s}}{\zeta_{X}(s)}=\lim_{s\to d^{+}}\frac{\zeta_{X}(s)-\zeta_{Z}(s)}{\zeta_{X}(s)}=1.$
We next address the crystalline compatibility. The proof uses the same principle as the proof of Theorem 4.4, except we need the existence of $p$$-\ell$ companions in our setting, as explained in [9]. By [19, Theorem 7.1], for any place $v|p$ of $E_{K_{0},s}$ not dividing $N$ the $p$-adic local system $\rho_{p,v}\colon\pi_{1}(\mathcal{S}_{(E_{K_{0},s})_{v}})\to G(\mathbb{Q}_{p})$ is crystalline. Our joint work [9] explains how to extract from this an ($G$-) overconvergent $F$-isocrystal, packaged as a $G(\overline{\mathbb{Q}}_{p})$-conjugacy class of representations
$\rho_{p,v}^{\operatorname{F-Isoc}^{\dagger}}\colon\pi_{1}^{\operatorname{F-Isoc}^{\dagger}}(\mathcal{S}_{v})\to G(\overline{\mathbb{Q}}_{p})$
compatible (via $D_{\mathrm{cris}}$) with $\rho_{p,v}$. For notational consistency, we will also for a place $\lambda$ of $\overline{\mathbb{Q}}$ above $p$ write these representations as $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda,v}$. We further know ([9, Proposition 4.3.1]) that the image of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ in $G^{\mathrm{ad}}$ is Zariski-dense, and from Lemma 4.1 we see that for any closed point $x\in\mathcal{S}_{v}$ that lifts to a special point, the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$ agrees in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ with $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ ($\lambda^{\prime}$ not above $p$). [^11] Here and in what follows, for $\lambda^{\prime}|p$ there is also a $\lambda^{\prime}$-companion, but since $\rho_{p,v}$ has $\mathbb{Q}_{p}$-coefficients, the “$p$-to-$p$” companion construction is just the identity. This compatibility on special points also implies the compatibility of the abelianizations of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda^{\prime},v}$. Combining work of Drinfeld ([7]), Abe ([1]), Abe-Esnault ([2]), and Kedlaya ([10]), $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ admits a $\lambda^{\prime}$-companion $\rho_{\lambda\rightsquigarrow\lambda^{\prime},v}$ as explained in [9, §4] whose algebraic monodromy group also contains $G^{\mathrm{der}}$.
Corollary 4.6.
For any place $v|p$ not above $N$ of $E^{\prime}$, and all finite places $\lambda|p$ and $\lambda^{\prime}$ of $\overline{\mathbb{Q}}$, $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda^{\prime},v}$ ($\lambda^{\prime}$ not above $p$) are $G$-companions.
Proof.
The adjoint case is the main result of [9]. By the observations in the paragraph preceding the Corollary (existence of the crystalline-étale companions on $\mathcal{S}_{v}$ and the previously-known, by Lemma 4.1, compatibility of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda^{\prime},v}$ on special points), the proof of Theorem 4.4 now applies verbatim. ∎
We conclude with a refinement of Corollary 4.6 that supplies one of the desiderata in the production of Kottwitz-triples (see the introduction). Let $v|p$ be a place not above $N$ of $E^{\prime}$, and let $x\in\mathcal{S}_{v}$ be a closed point. By construction, for all $\lambda^{\prime}$ not above $p$, $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$ lies in $G(\mathbb{Q}_{\ell^{\prime}})$, is semisimple by [3, Theorem 7.1], and we have seen this is an independent of $\lambda^{\prime}$ conjugacy class in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. For $\lambda|p$, let us write $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ for the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$. Corollary 4.6 tells us the semisimple part $\gamma_{x}$ of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ is $G(\overline{\mathbb{Q}}_{p})$-conjugate to an element of $G(\overline{\mathbb{Q}})$ whose conjugacy class is $\mathbb{Q}$-rational (induced by $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$). Note that we only at present know that $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$ is semisimple for $x$ in the $\mu$-ordinary locus, since $x$ then lifts to a special point in characteristic zero.
Corollary 4.7.
Let $K_{x}=\mathrm{Frac}(W(\kappa(x)))$. There is an element $\delta\in G(K_{x})$, defined up to $\sigma$-conjugacy by elements of $G(W(\kappa(x)))$, such that the $G(\overline{\mathbb{Q}}_{p})$-conjugacy class of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})$ equals that of $\delta\sigma(\delta)\cdots\sigma^{d_{x}-1}(\delta)$, where $d_{x}=[\kappa(x):\mathbb{F}_{p}]$.
Proof.
Recall that $K_{p}=G_{\mathbb{Z}_{p}}(\mathbb{Z}_{p})$ is a hyperspecial level structure, where we write $G_{\mathbb{Z}_{p}}$ for an extension of $G$ to a reductive model over $\mathbb{Z}_{p}$. Choose a faithful representation (over $\mathbb{Z}_{p}$) $L$ of $G_{\mathbb{Z}_{p}}$ and set $V=L\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Any choice of lift $\tilde{x}\in\mathcal{S}(W(\kappa(x)))$ induces on $L$ the structure of a crystalline $\pi_{1}(K_{x})$-representation, with underlying $F$-isocrystal independent of the choice of lift $\tilde{x}$. By [12, Proposition 1.3.2], there is a finite set of tensors $(s_{\alpha})\subset L^{\otimes}$ (the direct sum of all $\mathbb{Z}_{p}$-modules formed from $L$ using duals, tensor products, and symmetric and exterior powers) such that $G_{\mathbb{Z}_{p}}$ is the subgroup of $\mathrm{GL}(L)$ fixing (pointwise) all $s_{\alpha}$. Let $\mathfrak{M}$ be Kisin’s fully faithful tensor functor ([11], [12, §1.2])
$\mathfrak{M}\colon\mathrm{Rep}_{\pi_{1}(K_{x})}^{\mathrm{cris},\circ}\to\mathrm{Mod}^{\varphi}_{/\mathfrak{S}}$
from lattices in crystalline representations of $\pi_{1}(K_{x})$ [^12] Note Kisin’s theory works with ramified ground fields as well. to finite free $\mathfrak{S}=W(\kappa(x))[[u]]$-modules equipped with an isomorphism $\varphi^{*}(\mathfrak{M})[1/E(u)]\xrightarrow{\sim}\mathfrak{M}[1/E(u)]$; here $E(u)$ is an Eisenstein polynomial defining a fixed uniformizer of $K_{x}$, so we can just take $E(u)=u-p$.
The tensors $s_{\alpha}\in L^{\otimes}$ transport via $\mathfrak{M}$ to tensors $(\tilde{s}_{\alpha})\subset\mathfrak{M}(L)^{\otimes}$; Galois-invariance of the $s_{\alpha}$ yields $\varphi$-invariance of the $\tilde{s}_{\alpha}$, and then [12, Corollary 1.3.5] shows that the $(\tilde{s}_{\alpha})$ define a reductive group $G_{\mathfrak{M}}\subset\mathrm{GL}(\mathfrak{M}(L))$ isomorphic to $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$. More precisely, the isomorphism arises from trivializing the $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$-torsor $\mathcal{P}:=\underline{\mathrm{Isom}}_{\mathfrak{S}}(L\otimes_{\mathbb{Z}_{p}}\mathfrak{S},\mathfrak{M}(L))$ of linear isomorphisms mapping $s_{\alpha}$ to $\tilde{s}_{\alpha}$. Specializing along $u\mapsto 0$, we obtain tensors $\tilde{s}_{\alpha,0}\in(\mathfrak{M}(L)/u\mathfrak{M}(L))^{\otimes}$ and a corresponding trivialization.
Thus for a choice $\gamma$ of trivialization over $\mathfrak{S}/u\mathfrak{S}=W(\kappa(x))$, we have the $\sigma$-linear injection
[…]
that carries $s_{\alpha}$ to $s_{\alpha}$, since the tensors $\tilde{s}_{\alpha}$ and thus $\tilde{s}_{\alpha,0}$ are $\varphi$-invariant. The composite $\Phi$ above can therefore be uniquely expressed as $b\sigma$ for some $b\in G(K_{x})$. Changing the choice of trivialization of $\mathcal{P}\pmod{u}$ amounts to replacing $\gamma$ by $\gamma s$, $s\in G(W(\kappa(x)))$, which changes $\Phi$ to $s^{-1}\Phi s$ and changes $b$ to $s^{-1}b\sigma(s)$. Finally, $\Phi^{d_{x}}=b\sigma(b)\cdots\sigma^{d_{x}-1}(b)$ is an element of $G(K_{x})$ in the same $G(\overline{\mathbb{Q}}_{p})$-conjugacy class as $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$, since $\mathfrak{M}(L)/u\mathfrak{M}(L)[1/p]$ is canonically $D_{\mathrm{cris}}(V)$ as $\varphi$-isocrystals ([12, Theorem 1.2.1]). ∎